What was the grade a student received on her first examination, if her other grades were 70, 80, and 90, and her average on the four examinations was 75?
(A) 60 (B) 75 (C) 85 (D) 90 (E) 100

Answer by palanisamy(496)

The average grades received on the four examinations was 75.
Therefore total grades received on the four examinations = 4*(75) = 300
Her other grades were 70, 80, and 90
Therefore total grades received on the three examinations = 70+80+90 = 240
So the grade received in the first examination = 300-240
= 60